OpenStudy (anonymous):

What is the center and radius of the circle with the equation x2 + y2 + 8x – 4y – 61 = 0?

OpenStudy (anonymous):

Center: (–4, –2); Radius = 9 Center: (–4, 2); Radius = 9 Center: (4, –2); Radius = 9 Center: (4, 2); Radius = 9

OpenStudy (anonymous):

I think its the 2nd choice.

OpenStudy (anonymous):

Okie dokie. Thankyou(:

OpenStudy (anonymous):

\[\large x^2 + y^2 + 8x – 4y – 61 = 0 \] \[\large x^2 + 8x + y^2 – 4y – 61 = 0 \] \[\large (x^2 + 8x + 16) + (y^2 – 4y + 4) – 61-16-4 = 0 \] \[\large (x^2 + 8x + 16) + (y^2 – 4y + 4) – 81 = 0 \] \[\large (x+4)^2 + (y-2)^2 – 81 = 0 \] \[\large (x+4)^2 + (y-2)^2 = 81 \] \[\large (x+4)^2 + (y-2)^2 = 9^2 \] center is (-4,2) with radius 9